### v1

# Dom Huh    Prasant Mohapatra

Abstract
This paper proposes a diffusion-based auto-bidding framework that leverages graph representations to model large-scale auction environments. In such settings, agents must dynamically optimize bidding strategies under constraints defined by key performance indicator (KPI) metrics, all while operating in competitive environments characterized by uncertain, sparse, and stochastic variables. To address these challenges, we introduce a novel approach combining learnable graph-based embeddings with a planning-based latent diffusion model (LDM). By capturing patterns and nuances underlying the interdependence of impression opportunities and the multi-agent dynamics of the auction environment, the graph representation enable expressive computations regarding auto-bidding outcomes. With reward alignment techniques, the LDM’s posterior is fine-tuned to generate auto-bidding trajectories that maximize KPI metrics while satisfying constraint thresholds. Empirical evaluations on both real-world and synthetic auction environments demonstrate significant improvements in auto-bidding performance across multiple common KPI metrics, as well as accuracy in forecasting auction outcomes.
®ate Optimization
1 Introduction
Optimal auto-bidding in large-scale auctions remains an open research challenge with significant real-world implications for online markets. In these environments, auto-bidding agents must compete with bids on high-volume, stochastic streams of impression opportunities (IOs) to maximize value while adhering to multiple key performance indicator (KPI) metric constraints (Aggarwal et al., 2024; Kitts et al., 2017). Recent efforts (Guo et al., 2024) have demonstrated impressive successes with using generative diffusion models (Sohl-Dickstein et al., 2015) to simulate auction trajectories and plan bids to maximize their KPI adherence. However, existing methods rely on heuristic feature representations that inadequately capture the interdependencies between IOs and the multi-agent dynamics of auction environments.
Figure 1: An illustration of the LGD-AB framework. At every time-step, a graph is formed for agent $i$. The nodes of a sub-graph $\mathcal{G}_{t}^{i}$ represent the IO of agent $i$. Sub-graphs for other agents are similarly formed and connected accordingly. The graph is processed using a GNN to generate an embedding vector $x^{i}_{t}$. The diffusion model then learns a posterior of the sequence of embedding vectors.
To address this discrepancy, we introduce a graph-based representation of auto-bidding dynamics, modeling IOs on a more granular level to retain their individual attributes as well as encoding their connections to other IOs and agents. We then extend latent diffusion models (LDMs) (Rombach et al., 2022) to leverage this structure, formulating our framework – denoted Latent Graph Diffusion Model for Auto-Bidding (LGD-AB) – to forecast auction outcomes and extract crucial bidding information from a more expressive feature space. Further, we align the LDM (Huh & Mohapatra, 2025) to optimize against the KPI metrics. Our empirical experiments on real-world and synthetic auction simulations (Su et al., 2025; Deng et al., 2024) demonstrate that LGD-AB outperforms existing methods across common KPIs (He et al., 2021).
2 Latent Graph Diffusion Model for Auto-Bidding
Our proposed framework, LGD-AB, comprises two core components: (1) a graph-based embedding module for encoding auction dynamics and (2) a multi-agent latent diffusion model for auction forecasting and strategic bid optimization.
2.1 Graph Embedding For Ad Auctions
The objective of the graph embedding is to represent the IOs presented to each agent in an expressive and comprehensive manner. In large-scale auctions, the number of IOs exposed to each agent can range drastically, from 00 to $10$s of thousands at every time-step. To process data with such variability yet interdependency, we represent these IOs as nodes of a graph and process their features using graph convolutions with an attention-based graph neural network (GNN) (Brody et al., 2022). For scalability and computational limitations for larger scale simulations, we cluster IOs to avoid prohibitively large graphs. We additionally allocate a node dedicated for non-exposed IOs to maintain context during sparse time periods. In this graph representation, we fully connect the nodes belonging to same agents.
We optimize the latent graph embedding $x^{i}_{t}$ of agent $i$ at time-step $t$ generated by the GNN to hold information regarding the bids $b_{i}^{t}[k]$ made for each IO $k$. Hence, we introduce a separate inverse dynamics model (IDM) to support the optimization of the following objective:
$\mathcal{L}_{\text{graph}}(\theta)=\mathbb{E}_{(t,k)}[\log\text{P}(b_{t}^{i}[k]|\theta,x_{t}^{i},f_{t}^{i}[k],c_{t}^{i})]$ (1)
where $k$ is the IO identification index and $f_{t}^{i}[k]$ are the node features of the corresponding IO. To ensure the feasibility of this optimization, we append a context vector $c_{t}^{i}$ that provides more information regarding the agent agnostic to the individual IO. To clarify, we optimize the objective over the parameters $\theta$ of both the GNN and the IDM. We view this optimization to (1) instill bid-policy information into the embedding and (2) constrain the graph embedding to reflect auction dynamics.
Graphs of Other Agents
To ensure the embedding space accounts for the presence of other agents within the auction setting, we define additional sub-graphs representing the other agents constructed identically. The nodes between agents that share IOs are connected to encode the relational information that the exposures of IO can be shared between agents. IO nodes that are not exposed to an agent are connected to the unique IO node dedicated to non-exposed IOs, as seen in Figure 1.
Figure 2: Agent $i$’s approximation of Other Agents’ Graph $\omega^{i}(\mathcal{G}_{t}^{-i})$, with two nodes for student GNN.
Incomplete Information
We note that the advertiser’s visibility of other agents’ IOs and bids is contingent to the auction design and prevents the full construction of the graphs of other agents. In such cases, we extend the concept of centralized training and decentralized execution paired with knowledge distillation (KD) (Gou et al., 2021). We train a centralized GNN $\mathcal{T}$ with full visibility of all agents’ sub-graphs on Equation 1. We then train a student model $\mathcal{S}$ using a response-based KD loss to mimic $\mathcal{T}$. As shown in Figure 2, $\mathcal{S}$ approximates the other agents’ sub-graph with two proxy nodes with learnable node feature embeddings. Similarly, if the non-exposed IOs are not visible to the agent, we learn its node feature embedding.
Multi-Agent Latent State Optimization
We extend representation learning techniques, namely MA-SPL (Huh & Mohapatra, 2024), to further embed multi-agent and temporal dynamics information within the latent graph embedding $\{x^{i}_{0},x_{1}^{i},\dots,x_{T}^{i}\}_{i\in A}$, where $A$ is the set of all agents.
Approximate Equilibrium Computation
To promote a more cohesive multi-agent bidding strategy, we compute a joint embedding space by merging all agents’ graph embeddings and we refer to this design choice as approximate equilibrium computation (EC). Therefore, we introduce an aggregation operation $\mathcal{A}:\{x^{i}_{t}\}_{i\in A}\mapsto x_{t}$ using a self-attention network. This latent embedding $x_{t}$ is now shared between all agents when computing the embedding and generative objectives as well as for individual bid formulation.
2.2 Latent Diffusion Models
The latent diffusion model (LDM) processes a temporal sequence of graph embeddings $\{x_{t}\}_{t\in[\tau,\tau+T]}$, where $T$ is the length of the context window, and forms the foundation to devise a planning-based auto-bidding solution. In this work, we adopt the Decision Diffuser framework (Ajay et al., 2023), denoising state trajectories and offloading the action generation to the IDM. Unlike recent works using transformer-based architectures (Peebles & Xie, 2023) and score-based prediction (Song et al., 2020), we found greater success using the traditional 1D temporal convolution network with $\epsilon$-prediction. We optimize the diffusion objective using in-painting masks (Lugmayr et al., 2022).
Reward Alignment
To ensure that the LDM’s posterior aligns with the auto-bidding objective, we follow the procedure set in (Huh & Mohapatra, 2025), utilizing an iterative sequence of constrained alignment optimization of reinforcement learning, direct preference optimization, and supervised fine-tuning with rejection sampling. We train the value function using IQL (Kostrikov et al., 2021) to predict not only the return, i.e., the number of total conversations, but also other important KPI metrics, such as CPA, ROI, CVR, budget adherence, and social welfare. We define the conversion rate, i.e., CVR, as the ratio between the number of winning bids and the number of total bids made. With this value function, we can strictly enforce KPI constraints, most notably, budget adherence, through rejection sampling.
Fine-Tuning
Training LGD-AB occurred in two stages. We first trained the graph embedding module and then the LDM separately. We then train both modules together, fine-tuning the GNN at a lowered learning rate. We found that this additional stage of joint training benefits the performance of both modules.
3 Results and Ablations Studies
3.1 Experiment Setup
We evaluate LGD-AB on two auto-bidding simulations: AuctionNet (Su et al., 2025) and synthetic auction (Deng et al., 2024).
AuctionNet (Su et al., 2025) contains over $500$ million in real-world records of online advertising instances, with $21$ separate bidding periods of $48$ time-steps and $48$ different advertisers. We constructed an auto-bidding simulator using this dataset, following prior works (Chiappa et al., 2024).
Synthetic auction, on the other hand, describes a sound procedure to generate artificial auction data that mimics real ad auctions by capturing its underlying properties through hierarchical feature generation and specific probabilistic distributions. We extend (Deng et al., 2024) with configurable parameters to introduce longer and variable ad life cycles, multi-slot winning IOs, and a budget constraint on each agent. We collect an offline RL dataset of $100$ thousand auction simulations with a varying number of advertisers, allocation rule (e.g., FPA, GSP, VCG), bidding period durations, and IO’s properties. Within this dataset, advertisers either follow a uniform bid-scaling strategy or the non-uniform bid-scaling algorithm, as described in (Deng et al., 2024).
3.2 Empirical Evaluation
Auction Forecasting
We measure the forecasting capabilities of our LGD-AB framework by comparing the ground truth outcomes $x_{t^{\prime}>t}$ with the predictions of the LDM $z_{t^{\prime}>t}$ given a masked input $x_{t^{\prime}\leq t}$. To quantify this measure $F(\theta)$, we compute the log-likelihood of the LDM’s posterior to the path towards ground truth outcomes.
$F(\theta)=\mathbb{E}_{x,t}[\log\text{P}(x_{t^{\prime}>t}|\theta,x_{t^{\prime}\leq t})]$ (2)
In Table 1, our evaluation was performed using our online simulators, where to compute $F(\theta)$, we use a testing dataset that was not used during training. We found, with KD, the student model retains on average $77.48\%$ of the teacher model’s forecasting accuracy, despite partial observability. Interestingly, EC improves forecasting capability in Synthetic Auction, however, in AuctionNet, it failed to replicated this improvement. We attribute this to the limited number of data samples (i.e., only $21$ bidding periods in total), we utilize synthetic data in AuctionNet to mitigate the data scarity. After further training, we found a more fair representation of its forecasting accuracy. Lastly, the KPI alignment reduced average forecasting accuracy by $9.47\%$, which is expected, as such fine-tuning may lower the likelihood of lower value auto-bidding outcomes.
$F(\theta)$
AuctionNet Synthetic Auction
$\mathcal{T}$ $3.43$ $2.59$
$\mathcal{S}$ $2.64$ $2.02$
No EC $3.37$ $2.13$
EC $0.21\rightarrow 3.32^{*}$ $2.62$
No Alignment $3.25$ $2.09$
Alignment $3.02$ $2.04$
Table 1: Auction Forecasting Results.
AuctionNet Synthetic Auction
DiffBid LGD-AB DiffBid LGD-AB
Return $353.19\pm 0.42$ $489.04\pm 1.3$ $982.59\pm 10.61$ $2123.10\pm 15.32$
CPA $1.041\pm 0.028$ $0.992\pm 0.009$ $0.842\pm 0.014$ $0.592\pm 0.010$
ROI $0.012\pm 0.001$ $0.030\pm 0.001$ $0.026\pm 0.001$ $0.055\pm 0.001$
CVR $0.071\pm 0.002$ $0.152\pm 0.008$ $0.089\pm 0.001$ $0.203\pm 0.003$
Budget Adherence $0.995\pm 0.002$ $0.995\pm 0.002$ $0.998\pm 0.001$ $1.000\pm 0.000$
Social Welfare $3.63\pm 0.81$ $4.09\pm 0.07$ $50.43\pm 1.42$ $69.20\pm 0.56$
Table 2: Mean and ±$\pm 1$ std performances between DiffBid and LGD-AB ($\mathcal{S}$, EC, and post-alignment) over $64$ random seeds.
Figure 3: KPI Alignment Learning Curve on Synthetic Auction with multi-KPI alignment compared to optimizing each KPI criterion independently. For social welfare, we compare the training between EC and non-EC variants.
KPI and Performance
We benchmark our LGD-AB to a competitive diffusion-based planning solution, DiffBid (Guo et al., 2024) across six common KPI metrics. We highlight that the main differences of the DiffBid framework to our LGD-AB are its use of a heuristic feature space, bid-scaling, and a conditional return input. Comparisons to other generative models (i.e., Decision Transformer) and RL baselines (i.e., BC, CQL, and IQL) to DiffBid are provided in (Guo et al., 2024), where DiffBid has demonstrated notable improvements.
From Table 2, we observe improvements across all KPIs using our LGD-AB framework. The results demonstrate a significant increase in the number of IOs won, shown by average ×$\times 1.29$ improvement in return from our competitive baseline while significantly maximizing CVR and lowering CPA under adherence to the budget constraint. Lastly, our proposed LGD-AB has demonstrated better social welfare, indicating a collective utility to the auction. In both our baseline and LGD-AB solution, budget adherence was not an issue and remained comparable, maintaining near-perfect budget adherence over all trials.
As shown in Figure 3, we find that optimizing over multiple KPI (i.e., return, CPA, ROI, CVR, and social welfare) at the same time provides meaningful improvements within the individual performances of each metric. Additionally, the results affirm the utility of EC, as having a shared latent graph embedding can enable proper and more effective optimization of social welfare.
Accuracy of Bid Computation
We compare the bid prediction accuracy of DiffBid and LGD-AB in Table 3, depicting an improvement in being able to extract bids more accurately from our graph embeddings. Notably, in the Synthetic Auction simulation dataset, where portions of the auction data are generated from non-uniform bid-scaling strategies, we notice a wider margin of difference in accuracy in favor of our LGD-AB framework.
Average $\ell_{2}$ Norm per Agent
AuctionNet Synthetic Auction
DiffBid $19.78$ $82.10$
LGD-AB $13.41$ $37.82$
Table 3: Comparsion between Bid Computation Accuracy of DiffBid and LGD-AB.
Limitations and Future Directions
We observe two key issues and limitations while developing our proposed solution, mainly pertaining to the scale of data and efficiency of our generative modeling algorithm. With AuctionNet, while there exists a large set of individualized auction records to train on, there was an insufficient number of collective auction records that span full bidding periods to properly train a generative model to capture the underlying auction dynamics accurately. We addressed this limitation with synthetic data generation and hyper-parameter tuning on the diffusion model. However, this limited performance can be improved with advancements in either training generative models in low data regimes (e.g., few-shot learning) or greater data availability. The other issue concerns the computational limitation from scaling the graph representation w.r.t. the number of IOs exposed to each agent. In theory, the original proposition was to allocate a node for each IO with edges denoting various aspects of their pairwise inter-relations, which scales poorly with a large number of shared IOs; an important feature was noticed in AuctionNet. Again, while we managed to address this with a clustering approach for AuctionNet, this reduced representation significantly reduced the expressiveness of the graph embeddings. In future works, a form of dynamic graph sparsification or hierarchical graph representations could mitigate this.
4 Conclusion
In this work, we propose a latent diffusion planning framework using graph-based embedding for auto-bidding applications. We demonstrate the utility of a graph-based representation, which enabled the LDM to capture the underlying dynamics distribution of a multi-agent auction process. Through aligning various KPIs on the trained LDM, we devised an effective bidding strategy capable of generating bids that maximize various KPI metrics.